# Magnetar structure in nonlinear electrodynamics with mixed poloidal-toroidal fields

### v1

@@ 5.4 Crustal failures
Figure 6: Dimensionless ratios of BI-to-Maxwell stresses (48) in the $r\theta$ (left) and $\theta\theta-\phi\phi$ (right) directions within a ‘crust’ for the field (27), with redder shades indicating a greater ratio $\mathcal{R}$. The classical polar strength is fixed as $B_{p}=b/2$.
@@ 6 Discussion
We also explore how limitations on the toroidal partition together with the effective augmentation of the near-surface field impacts on magnetar phenomena. For example, we find that because the current is effectively reduced in NLED via the generalised Ampére law (1), the effective Ohmic and Hall timescales are reduced slightly for super-Schwinger fields (see Fig. 5). This may help to explain the coldness and longevity of strong fields in the mysterious class of ULPs, assuming (a subset of them) are isolated magnetars (Hurley-Walker et al., 2022; Suvorov & Melatos, 2023; de Ruiter et al., 2024), especially since spindown is also accelerated for strong fields (see equation 47). A thorough investigation of Hall-Ohm (and magnetothermal more generally) evolution in NLED, including crustal plasticity (Sec. 5.4), is reserved for future work.
Several extensions of this study are natural. One path involves the inclusion of general-relativistic effects, which may have interesting interplays with NLED as the energy density $T^{tt}$ becomes large (Rezzolla & Ahmedov, 2004; Pétri, 2015). It has been argued, for instance, that the Blandford-Znajek process is less effective in NLED (Li & Wang, 2020), which could influence the properties of relativistic jets relevant for GRBs (Sec. 5.3). Explicit solutions to the GGS equation, either for non-barotropic or barotropic stars (compare the calculations in Secs. 3.2 and Appendix A), could also be useful to study the stability of internal fields along the lines discussed by Rau & Wasserman (2021) for the core or Rau & Wasserman (2023) for the crust. It would also be worth tackling the magnetosphere problem in more detail. Using the methodology introduced by Urbán et al. (2023) and Stefanou, Urbán & Pons (2023), the force-free NLED structure of the magnetosphere could be solved for, including the possibility of rotation (‘NLED pulsar equation’). The structure of coronal loops, for example, relating magnetic topology to flare properties and hotspots may differ because the effective twist is modulated by $L^{\prime}(F^{2})$, as per equation (7).
@@ Acknowledgements
Support provided by the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana through Prometeo Project CIPROM/2022/13 is acknowledged. AGS extends his thanks to the members of the theoretical astrophysics group in Tübingen for the warm welcome he received while this work was being completed.
@@
Appendix A Hydromagnetic equilibria for constant-density stars
@@
where
$\tilde{\Delta}\psi=\frac{\partial}{\partial r}\left[L^{\prime}\left(\frac{\nabla\psi\cdot\nabla\psi}{2r^{2}\sin^{2}\theta}\right)\frac{\partial\psi}{\partial r}\right]+\frac{\sin\theta}{r^{2}}\frac{\partial}{\partial\theta}\left[\frac{L^{\prime}\left(\frac{\nabla\psi\cdot\nabla\psi}{2r^{2}\sin^{2}\theta}\right)}{\sin\theta}\frac{\partial\psi}{\partial\theta}\right].$ (51)
It is clear that the set of magnetostatic equilibria in this case (and barotropic cases more generally) are restricted because the field must now satisfy equation (50); see Glampedakis & Lasky (2016) for a discussion on the severity of this constraint.

### v2

@@ 5.4 Crustal failures
Figure 6: Dimensionless ratios of BI-to-Maxwell stresses (48) in the $r\theta$ (left) and $\theta\theta-\phi\phi$ (right) directions within a ‘crust’ for the field (3.2), with redder shades indicating a greater ratio $\mathcal{R}$. The classical polar strength is fixed as $B_{p}=b/2$.
@@ 6 Discussion
We also explore how limitations on the toroidal partition together with the effective augmentation of the near-surface field impacts on magnetar phenomena. For example, we find that because the current is effectively reduced in NLED via the generalised Ampére law (1), the effective Ohmic and Hall timescales are reduced slightly for super-Schwinger fields (see Fig. 5). This may help to explain the coldness and longevity of strong fields in the mysterious class of ULPs, assuming (a subset of them) are isolated magnetars (Hurley-Walker et al., 2022; Suvorov & Melatos, 2023; Cooper & Wadiasingh, 2024), especially since spindown is also accelerated for strong fields (see equation 47). A thorough investigation of Hall-Ohm (and magnetothermal more generally) evolution in NLED, including crustal plasticity (Sec. 5.4), is reserved for future work.
Several extensions of this study are natural. One path involves the inclusion of general-relativistic effects, which may have interesting interplays with NLED as the energy density $T^{tt}$ becomes large (Rezzolla & Ahmedov, 2004; Pétri, 2015). It has been argued, for instance, that the Blandford-Znajek process is less effective in NLED (Li & Wang, 2020), which could influence the properties of relativistic jets relevant for GRBs (Sec. 5.3). Explicit solutions to the GGS equation, either for non-barotropic or barotropic stars (compare the calculations in Secs. 3.2 and Appendix B), could also be useful to study the stability of internal fields along the lines discussed by Rau & Wasserman (2021) for the core or Rau & Wasserman (2023) for the crust. It would also be worth tackling the magnetosphere problem in more detail. Using the methodology introduced by Urbán et al. (2023) and Stefanou, Urbán & Pons (2023), the force-free NLED structure of the magnetosphere could be solved for, including the possibility of rotation (‘NLED pulsar equation’). The structure of coronal loops, for example, relating magnetic topology to flare properties and hotspots may differ because the effective twist is modulated by $L^{\prime}(F^{2})$, as per equation (7).
@@ Acknowledgements
We thank the anonymous referee for constructive comments. Support provided by the Conselleria d’Educació, Cultura, Universitats i Ocupació de la Generalitat Valenciana through Prometeo Project CIPROM/2022/13 is acknowledged. AGS extends his thanks to the members of the theoretical astrophysics group in Tübingen for the warm welcome he received while this work was being completed.
@@
Appendix A Radial prefactors in magnetic field expressions
In an effort to preserve the flow of text, details regarding derivations of a few lengthy expressions are presented here rather than the main body.
Consider the force-free equation (17), $0=\tilde{\Delta}\psi$, where
$\tilde{\Delta}\psi=$ $\frac{\partial}{\partial r}\left[L^{\prime}\left(\frac{\nabla\psi\cdot\nabla\psi}{2r^{2}\sin^{2}\theta}\right)\frac{\partial\psi}{\partial r}\right]$ (50)
$+\frac{\sin\theta}{r^{2}}\frac{\partial}{\partial\theta}\left[\frac{L^{\prime}\left(\frac{\nabla\psi\cdot\nabla\psi}{2r^{2}\sin^{2}\theta}\right)}{\sin\theta}\frac{\partial\psi}{\partial\theta}\right].$
Using the PPM Lagrangian function (15), $L(F^{2})=-F^{2}+\alpha_{\rm PM}F^{4}+\alpha_{\rm PPM}F^{6}$, an approximate dipolar solution, $\psi(r,\theta)\approx f(r)\sin^{2}\theta$, can be found using a Legendre projection. Explicitly, inserting this form of $\psi$ into equation (50), multiplying the result by $dP_{1}(\cos\theta)/d\theta=\cos\theta$, and then integrating over angles ($0\leq\theta\leq\pi$) leads to the nonlinear, ordinary differential equation (omitting the implied radial dependence)
$0=$ $\left(\frac{2f}{r^{2}}-f^{\prime\prime}\right)+\frac{4}{5}\alpha_{\rm PM}\Big{[}\frac{3(f^{\prime})^{2}f^{\prime\prime}}{r^{2}}-\frac{2(f^{\prime})^{3}}{r^{3}}$ (51)
$+\frac{f(f^{\prime})^{2}+f^{2}f^{\prime\prime}}{r^{4}}-\frac{4f^{2}f^{\prime}}{r^{5}}-\frac{6f^{3}}{r^{6}}\Big{]}$
$+\frac{18}{35}\alpha_{\rm PPM}\Big{[}\frac{5(f^{\prime})^{4}f^{\prime\prime}}{r^{4}}-\frac{4(f^{\prime})^{5}}{r^{5}}$
$+\frac{2f(f^{\prime})^{4}+4f^{2}(f^{\prime})^{2}f^{\prime\prime}}{r^{6}}-\frac{8f^{2}(f^{\prime})^{3}}{r^{7}}$
$+\frac{4f^{3}(f^{\prime})^{2}+2f^{4}f^{\prime\prime}}{r^{8}}-\frac{16f^{4}f^{\prime}}{r^{9}}-\frac{20f^{5}}{r^{10}}\Big{]},$
which has the expected solution $f(r)=B_{p}/2r$ for $\alpha_{\rm PM}=\alpha_{\rm PPM}=0$. In general though, equation (51) resists transform and other integration techniques and we cannot find a general solution. Instead, we seek a series expansion of the form
$f(r)=\sum_{j=0}^{J}\sum_{k=0}^{K}f_{j,k}(r)(\alpha_{\rm PM})^{j}(\alpha_{\rm PPM})^{k},$ (52)
for some functions $f_{j,k}$ and integers $J,K$, taking care to ensure dimensional consistency. It turns out that the functions $f_{j,k}$ can be found exactly by matching orders. The case considered in the main text corresponds to $J=2$ and $K=0$ (expression 21), defining the second-order PM expansion. For instance, the inhomogenous $\mathcal{O}(\alpha_{\rm PM})$ equation (sourced by Maxwell terms with $f_{0,0}=B_{p}/2r$) reads
$0=\frac{9B_{p}^{3}}{10r^{9}}+\frac{2f_{1,0}(r)}{r^{2}}-f_{1,0}^{\prime\prime}(r),$ (53)
which has the particular solution
$f_{1,0}(r)={B_{p}^{3}}/{60r^{7}},$ (54)
which is precisely the second term appearing in expression (21) from the main text. Having found the solution (54), as in other works (e.g. Li, Yang & Wang, 2020), the $\mathcal{O}(\alpha_{\rm PM}^{2})$ equation (sourced by first-order corrections) becomes
$0=\frac{129B_{p}^{5}}{100r^{15}}+\frac{2f_{2,0}(r)}{r^{2}}-f_{2,0}^{\prime\prime}(r),$ (55)
which has the solution $f_{2,0}(r)=43B_{p}^{5}/6000r^{13}$. This matches the last term in expression (21). Note that these cumbersome numerical prefactors appear because of the factors of 2 in both the denominator within the argument of $L$ and the zeroth-order solution, together with the fact that we have opted not to normalise the PM parameters by factorials to avoid corrupting interpretations of the orders.
For the BI (or similar) theory where there is only one scale parameter, $b$, the double sum (52) collapses to
$f(r)=\sum_{j=0}^{J}f_{j}(r)b^{-2j},$ (56)
recalling that only even powers of $1/b$ appear in the PM expansion of the Born-Infeld action; see expression (12). Repeating the same exercise as above but in inverse powers of $b$ at one-higher order, $J=3$, one finds $f_{0}=B_{p}/2r,f_{2}=B_{p}^{3}/120r^{7}$ (as above, recalling that $\alpha_{\rm PM}=1/2b^{2}$) and
$0=\frac{93B_{p}^{4}}{2800r^{15}}+\frac{2f_{4}(r)}{r^{2}}-f_{4}^{\prime\prime}(r)=0,$ (57)
with solution $f_{4}(r)=31B_{p}^{5}/168000r^{13}$ and
$0=-\frac{16437B_{p}^{6}}{4r^{21}}+\frac{2f_{6}(r)}{r^{2}}-f_{6}^{\prime\prime}(r)=0,$ (58)
with solution $f_{6}(r)=-5479B_{p}^{7}/14112000r^{19}$. It is not difficult to recognise that the general solution takes the form $f_{j}(r)\propto B_{p}^{j+1}r^{-3j-1}$ with appropriate coefficients; the $\mathcal{O}(b^{-6})$ solution found above is that reported in expression (23). Though clumsy, this method can be extended to arbitrary values of $J,K$ (plus their extensions to PPPM and so on) and multipolarity to build perturbative, potential-field solutions in any given nonlinear theory of electrodynamics.
Deriving expression (3.2) from the main text for the radial prefactor of the internal field $(\psi_{I})$ follows the same order-matching logic. Through a slight abuse of notation we take
$\psi_{I}=f_{I}(r)\sin^{2}\theta=\sum_{i=2,4,6,8}\sum_{j=0,1,2,3}a_{ij}b^{-2j}r^{i}\sin^{2}\theta,$ (59)
for which we have sixteen free constants $a_{ij}$. We match the above at the surface $(r=1)$ to expression (56) describing the external solution $(\psi_{E})$ at each power of $b$: $\psi_{I}(1,\theta)-\psi_{E}(1,\theta)=0+\mathcal{O}(b^{-8})$, together with its first-order derivatives, $\left(\partial\psi_{I}/\partial r-\partial\psi_{E}/\partial r\right)|_{r=1}=0+\mathcal{O}(b^{-8})$. This constrains eight of the $a_{ij}$. The current, found from expression (1), $\bm{J}=-c\nabla\times\left[L^{\prime}(F^{2})\bm{B}\right]/4\pi$, can similarly be decomposed into inverse powers of $b$ (and/or multipoles). Noting that $\bm{J}$ is only azimuthal at the surface because we match to a potential field, the condition $\bm{J}_{\phi}|_{r=1}=0+\mathcal{O}(b^{-8})$ adds four constraints for a total of twelve. The matching problem thereby reduces to one of linear algebra, which can be straightforwardly handled with software-assisted Gaussian elimination. Though conceptually simple, the matrix equations are lengthy even for an Appendix. Still, carrying this out returns expression (3.2) from the main text, noting there remains four free constants (labelled $d_{i}$ for $i=0,1,2,3$ in Sec. 3.2) because we impose twelve conditions. This was done to explore a wider range of theoretical equilibria.
Appendix B Hydromagnetic equilibria for constant-density stars
@@
where $\tilde{\Delta}\psi$ is given in (50). It is clear that the set of magnetostatic equilibria in this case (and barotropic cases more generally) are restricted because the field must now satisfy equation (60); see Glampedakis & Lasky (2016) for a discussion on the severity of this constraint.